show work! thank you!Pearl Products Limited of Shenzhen, China, manufactures and distributes toys throughout South East Asia. Three cubic centimet3. Prepare a direct materials budget showing the quantity of solvent H300 to be purchased for July, August, and September, an
Answer #1
1
Pearl Products Limited
Production Budget
July August September October
Budgeted unit sales 37000 42000 52000 32000
Add: Desired units of ending finished goods inventory 12500 15000 10000 7500
Total needs 49500 57000 62000 39500
Less: Units of beginning finished goods inventory 11250 12500 15000 10000
Required production in units 38250 44500 47000 29500
3
Pearl Products Limited
Direct Materials Budget
July August September Third Quarter
Units of raw materials needed to meet production 114750 133500 141000 389250
Add: Desired units of ending raw materials inventory 66750 70500 44250 44250
Total units of raw materials needed 181500 204000 185250 433500
Less: Units of beginning raw materials inventory 56000 66750 70500 56000
Units of raw materials to be purchased 125500 137250 114750 377500
Workings:
Desired units of ending finished goods inventory
July = 2000+(42000*25%)= 12500
August = 2000+(52000*25%)= 15000
September = 2000+(32000*25%)= 10000
October = 2000+(22000*25%)= 7500
Units of raw materials needed to meet production:
July = 38250*3 = 114750
August = 44500*3 = 133500
September = 47000*3 = 141000
Desired units of ending raw materials inventory:
July = 133500*50% = 66750
August = 141000*50% = 70500
September = 29500*3*50%=44250